OpenStudy (anonymous):

I really want to know how to solve a the I, R, A of a circuit that contains a diode.

OpenStudy (kenljw):

The approximate diode current is Id = Is EXP(Vd/VT), where VT = KT/q = .026 V @T= 298 degrees K, Is is the saturation current, Vd is voltage across diode. 1/rd = d[Id}/dVd = Is EXP(Vd/VT) /VT = (Id/VT) You can be much more exact using exact equation for diode with DC and signal voltages across it which may depend on series resistance with it. This may become complicated, the analysis is usually done when the diode I used in AM demodulation circuit.

OpenStudy (anonymous):

thank you so much for the reply. I hope that you can make all the students love what they are doing. Let them know what makes passion to be as good as reaching your goals in life. God bless!